# Pass-efficient Randomized Algorithms for Low-rank Approximation of Quaternion Matrices

### v1

@@ 2 Preliminaries
[…]
@@
The conjugate of a quaternion $\mathbf{q}$ is defined as: $\mathbf{q}^{*}=q_{0}-q_{1}\mathbf{i}-q_{2}\mathbf{j}-q_{3}\mathbf{k}$. The norm (or magnitude) of a quaternion $\mathbf{q}$ is given by:$\|\mathbf{q}\|=\sqrt{q_{0}^{2}+q_{1}^{2}+q_{2}^{2}+q_{3}^{2}},$ and the inverse of a non-zero quaternion $\mathbf{q}$ is given by: $\mathbf{q}^{-1}=\frac{\mathbf{q}^{*}}{\|\mathbf{q}\|^{2}}.$
@@
A quaternion matrix is a matrix whose elements are quaternions. The space of quaternion matrices of size $m\times n$ is denoted by $\mathbb{Q}^{m\times n}$. The conjugate transpose of a quaternion matrix ${\bf X}={\bf X}_{1}+{\bf X}_{2}{\bf i}+{\bf X}_{3}{\bf j}+{\bf X}_{4}{\bf k}$ is defined as ${\bf X}^{H}={\bf X}^{T}_{1}-{\bf X}^{T}_{2}{\bf i}-{\bf X}^{T}_{3}{\bf j}-{\bf X}^{T}_{4}{\bf k}$, and its Frobenius norm is defined as $||{\bf X}||_{F}=\sqrt{\sum_{i,j}||x_{i,j}||^{2}}.$ A quaternion matrix ${\bf V}\in\mathbb{Q}^{n\times n}$ is said to be unitary if ${\bf V}^{H}{\bf V}={\bf I}$, where $V^{H}$ denotes the Hermitian of V.
@@ 3 Proposed pass-efficient algorithms for quaternion matrices
The approximation described above can be improved by using the oversampling concept, which means the sampling matrix ${\bf\Omega}\in\mathbb{Q}^{n\times(k+p)}$ instead of a sampling matrix of size $n\times k$. This helps to better capture the range of the matrix ${\bf X}$. However, this approach is applicable only when the singular values of the matrix ${\bf X}$ decrease quite rapidly. To handle the scenario of matrices with low decaying singular values, the power iteration scheme is used. Power iteration can be used to refine these approximations by improving the accuracy of the computed eigenvectors or singular vectors. It improves the alignment ${\bf Y}={({\bf X}{\bf X}^{H})^{q}{\bf X}{\bf\Omega}}$ with the dominant singular vectors, which $q$ is the number of power iterations. Increasing $q$ improves the accuracy of the approximation, while each power iteration requires additional matrix multiplications, increasing the computational cost. The randomized algorithm equipped with the oversampling and power iteration scheme is presented in Algorithm 1. An essential property of this algorithm is that for a given power iteration parameter $q$, the algorithm requires viewing or passing the data matrix ${\bf X},$ $22q+$ 2 times. More precisely, $2q$ times in lines 4-5 inside the “for” loop and two additional passes in lines 2 and 7. So, the question is whether it is possible to have a randomized algorithm that can provide a low-rank approximation for any budget of views, not necessarily even numbers. This question was answered positively in [19] and the author suggested a flexible algorithm in this sense, and it was later extended to tensors in [20]. Here, we use this technique for quaternion matrices. The algorithm is summarized in Algorithm 2. To be more precise, Algorithm 2, for a given budget $v\geq 2$ of views, computes an approximate truncated SVD $[{\bf U}_{R},{\bf\Sigma}_{R},{\bf V}_{R}]$ using one of the following strategies:
@@
    (If $v$ is even) Compute an orthonormal matrix ${\bf Q}^{(2)}$ whose columns build a basis for the column space of $({\bf X}{\bf X}^{H})^{(v-2)/2}{\bf X}{\bf\Omega}$. Then compute the rank-R TSVD $[{\bf V}_{R},{\bf\Sigma}_{R},\widetilde{{\bf U}_{R}}]$ of ${\bf X}^{H}{\bf Q}^{(2)}$ and set ${\bf U}_{R}={\bf Q}^{(2)}\widetilde{{\bf U}_{R}}$.
@@
    (If $v$ is odd) Compute an orthonormal matrix Qr whose columns form a basis for the column space of $({\bf X}^{H}{\bf X})^{(v-1)/2}{\bf\Omega}$. Then compute the rank-R TSVD $[{\bf U}_{R},{\bf\Sigma}_{R},\widetilde{{\bf V}_{R}}]$ of ${\bf X}{\bf Q}^{(1)}$ and set ${\bf V}_{R}={\bf Q}^{(1)}\widetilde{{\bf V}_{R}}$.
@@ 6       $[{\bf Q}^{(1)},\sim]={{\rm QR}}({\bf X}{\bf Q}^{(2)})$;
9$[{\bf Q}^{(2)},{\bf R}]={{\rm QR}}({\bf X}^{H}{\bf Q}^{(1)})$;
10 $[\widehat{\bf V},{\bf S},\widehat{\bf U}]={\rm Truncated}$ $\,{\rm SVD}({\bf R},k)$;
11 ${\bf V}={\bf Q}^{(2)}\widehat{\bf V}$;
${\bf U}={\bf Q}^{(1)}\widehat{\bf U}$;
@@ 4 Theoretical results
This section is devoted to computing an upper bound for the expected approximation error, i.e. $||\hat{{\bf X}}^{(q)}_{k+p}-{\bf X}||_{2}=||({\bf I}-{\bf Q}^{(1)}{\bf Q}^{(1)H}){\bf X}||_{2},\,q\geq 0,$ of the approximations computed by the proposed algorithm. The following lemma presents an upper bound for the expected approximation error (in spectral norm) for a rank-$(k+p)$ matrix of Algorithm 1.
@@
$E(||{\bf X}-{\bf Q}^{(1)}{\bf Q}^{(1)H}{\bf X}||_{2})\leq\Big{(}\big{(}1+3\sqrt{\frac{k}{4p+2}}\big{)}\sigma_{k+1}^{v}+\frac{3e\sqrt{4k+4p+2}}{2p+2}(\sum\limits_{j>k}\sigma_{j}^{2v})^{\frac{1}{2}}\Big{)}^{\frac{1}{v}}.$ (7)
$E(||{\bf X}-{\bf X}{\bf Q}^{(2)}{\bf Q}^{(2)H}||_{2})\leq\Big{(}\big{(}1+3\sqrt{\frac{k}{4p+2}}\big{)}\sigma_{k+1}^{v-1}+\frac{3e\sqrt{4k+4p+2}}{2p+2}(\sum\limits_{j>k}\sigma_{j}^{2(v-1)})^{\frac{1}{2}}\Big{)}^{\frac{1}{v-1}}.$
From ${\bf V}={\bf Q}^{(2)}\widehat{\bf V}$ and ${\bf U}={\bf Q}^{(1)}\widehat{\bf U}$, we deduce that the right and left singular vectors are as a linear combination of the columns ${\bf Q}^{(1)}$ and ${\bf Q}^{(2)}$, respectively. Moreover, relation (4) shows that for an even number of matrix views $v$, ${\bf V}$ can be more accurate than or as accurate as ${\bf U}$. If $v$ is an odd number, then relation (7) shows that ${\bf U}$ is more accurate than or as accurate as ${\bf V}$. Hence, for approximating the right singular vectors of matrix ${\bf X}$, when the number of views $v$ is an even number, it can be more accurate to apply Algorithm 2 to ${\bf X}$. Also, when $v$ is an odd number and our goal is to approximate the right singular vectors, we apply Algorithm 2 to ${\bf X}^{H}$.

### v2

@@ 2 Preliminaries
We denote matrices by capital bold letters, e.g. $\bf X$}. The spectral norm of matrices, the Euclidean norm of vectors, and the mathematical expectation are denoted by $\|.\|$, $\|.\|_{2}$ and $E$. A quaternion is typically represented as:
@@
[…]
@@
[…]
@@ 3 Proposed pass-efficient algorithms for quaternion matrices
The approximation described above can be improved by using the oversampling concept, which means the sampling matrix ${\bf\Omega}\in\mathbb{Q}^{n\times(k+p)}$ instead of a sampling matrix of size $n\times k$. This helps to better capture the range of the matrix $\bf X$}. However, this approach is applicable only when the singular values of the matrix $\bf X$} decrease quite rapidly. To handle the scenario of matrices with low decaying singular values, the power iteration scheme is used. Power iteration can be used to refine these approximations by improving the accuracy of the computed eigenvectors or singular vectors. It improves the alignment ${\bf Y}={({\bf X}{\bf X}^{H})^{q}{\bf X}{\bf\Omega}}$ with the dominant singular vectors, which $q$ is the number of power iterations. Increasing $q$ improves the accuracy of the approximation, while each power iteration requires additional matrix multiplications, increasing the computational cost. The randomized algorithm equipped with the oversampling and power iteration scheme is presented in Algorithm 1. An essential property of this algorithm is that for a given power iteration parameter $q$, the algorithm requires viewing or passing the data matrix $\bf X$}, $2q+2$ times. More precisely, $2q$ times in lines 4-5 inside the “for” loop and two additional passes in lines 2 and 7. So, the question is whether it is possible to have a randomized algorithm that can provide a low-rank approximation for any budget of views, not necessarily even numbers. This question was answered positively in [19] and the author suggested a flexible algorithm in this sense, and it was later extended to tensors in [20]. Here, we use this technique for quaternion matrices. The algorithm is summarized in Algorithm 2. To be more precise, Algorithm 2, for a given budget $v\geq 2$ of views, computes an approximate truncated SVD ${\bf U}_{R},{\bf\Sigma}_{R},{\bf V}_{R}]$ using one of the following strategies:
@@
    (If $v$ is even) Compute an orthonormal matrix ${\bf Q}^{(2)}$ whose columns build a basis for the column space of $({\bf X}{\bf X}^{H})^{(v-2)/2}{\bf X}{\bf\Omega}$. Then compute the rank-R TSVD ${\bf V}_{R},{\bf\Sigma}_{R},\widetilde{{\bf U}_{R}}$] of ${\bf X}^{H}{\bf Q}^{(2)}$ and set ${\bf U}_{R}={\bf Q}^{(2)}\widetilde{{\bf U}_{R}}$.
@@
    (If $v$ is odd) Compute an orthonormal matrix Qr whose columns form a basis for the column space of $({\bf X}^{H}{\bf X})^{(v-1)/2}{\bf\Omega}$. Then compute the rank-R TSVD ${\bf U}_{R},{\bf\Sigma}_{R},\widetilde{{\bf V}_{R}}$] of ${\bf X}{\bf Q}^{(1)}$ and set ${\bf V}_{R}={\bf Q}^{(1)}\widetilde{{\bf V}_{R}}$.
@@ 6  ${\bf Q}^{(1)},\sim]={{\rm QR}}({\bf X}{\bf Q}^{(2)})$;
[…]
10 $\widehat{\bf V},{\bf S},\widehat{\bf U}]={\rm Truncated}$ ${\rm SVD}({\bf R},k)$;
11 $^{\bf V}={\bf Q}^{(2)}\widehat{\bf V}$;
$^{\bf U}={\bf Q}^{(1)}\widehat{\bf U}$;
@@ 4 Theoretical results
[…]
@@
[…]
From $^{\bf V}={\bf Q}^{(2)}\widehat{\bf V}$ and $^{\bf U}={\bf Q}^{(1)}\widehat{\bf U}$, we deduce that the right and left singular vectors are as a linear combination of the columns ${\bf Q}^{(1)}$ and ${\bf Q}^{(2)}$, respectively. Moreover, relation (4) shows that for an even number of matrix views $v$, $\bf V$} can be more accurate than or as accurate as $\bf U$}. If $v$ is an odd number, then relation (7) shows that $\bf U$} is more accurate than or as accurate as $\bf V$}. Hence, for approximating the right singular vectors of matrix $\bf X$}, when the number of views $v$ is an even number, it can be more accurate to apply Algorithm 2 to $\bf X$}. Also, when $v$ is an odd number and our goal is to approximate the right singular vectors, we apply Algorithm 2 to ${\bf X}^{H}$.